Count the number of valence electron per metal of the following compounds and then, elucidate the structure.
(a) Ru3 (μ-CO)2 (CO)8 (bipy)
(b) RuCl (PPh3) (η1–Ph2PCH2PPh2) (η5–C5H5)
1
Expert's answer
2019-11-27T07:12:20-0500

(18 × n) – A, n is the number of metals in the complex

a) Ru3(μ-CO)2(CO)8

(18 × 3) - 42 = 54 - 42 = 12

b) RuCl(PPh3)(η1–Ph2PCH2PPh2)(η5–C5H5)

(18 × 1) - 16 = 2